OpenStudy (anonymous):

Integrate ln(1+x^2).

myininaya (myininaya):

my first thought is integration by parts

OpenStudy (anonymous):

I had this on one of my homework problems for calc 2. The answer the professor ultimately gave was different from what I had, and so was what wolfram alpha gave me. \[\int\limits_{ }^{} \ln \left( 1+x^2 \right)dx\] I set u=1+x^2 and thus had \[\int\limits_{ }^{} \ln \left(u \right)du\] Which can be easily integrated, but when you put u back in it isn't the same as what was on the answer sheet.

myininaya (myininaya):

that is because du doesn't equal dx there

OpenStudy (anonymous):

How do you integrate a natural log function by parts if there's no multiplication involved?

myininaya (myininaya):

times 1

myininaya (myininaya):

\[\int\limits \ln(1+x^2) dx=\int\limits 1 \cdot \ln(1+x^2) dx\]

OpenStudy (anonymous):

So, lnx=u and x=v?

myininaya (myininaya):

u=ln(1+x^2) and dv=1 dx

OpenStudy (anonymous):

just had the entire thing worked out and awaiting confirmation and I accidentally deleted it fml

myininaya (myininaya):

are you doing okay?

OpenStudy (anonymous):

\[u=\ln(1+x^2)\] \[dv=1dx\] \[v=x\] \[du=\frac{ 2x }{ x^2+1 }\] \[\int\limits_{}^{}(\ln(1+x^2)dx=(\ln(1+x^2)*(x))-\int\limits_{}^{}(x)(\frac{ 2x }{ x^2+1 })dx\] That last integral turns into a mess.

myininaya (myininaya):

\[\frac{2x^2}{x^2+1}=\frac{2(x^2+1)-2}{x^2+1}=2-\frac{2}{x^2+1}\]

OpenStudy (anonymous):

I see. Let me work this out on paper and I'll get back to you. Thanks for your help and your patience, I appreciate it.

myininaya (myininaya):

you might need to holler at someone else i have to go 2 should be easy to integrate integrating1/(x^2+1) should be from memory that this is arctan(x)

OpenStudy (anonymous):

That would explain the 2tan^-1(x) in the answer. Still haven't memorized that.